A sample of a breathing mixture for divers contained 34.3% helium, He; 51.7% nitrogen, N2; and 14.0% oxygen, O2 (by mass). What is the density of this mixture at 23°C and 772 mmHg?

d= g/L
Answer #1

34.3% He mean 34.3g of He present in 100g of gases

no of moles of He (nHe) = W/G.A.Wt

                                          = 34.3/4   = 8.575 moles

51.7% He mean 51.7g of N2 present in 100g of gases

no of moles of N2 (nN2)   = W/G.M.Wt

                                        = 51.7/28   = 1.85moles

14% O2 mean 14g of O2 present in 100g of gases

no of moles of O2( nO2)   = W/G.M.Wt

                                          = 14/32   = 0.4375 moles

Total no of moles of mixture (n) = nHe + nN2 + nO2

                                                    = 8.575 + 1.85 + 0.4375

                                                      = 10.8625moles

Total 100g of gases mixture

T = 23+273 = 296K

P = 722/760   = 0.95atm

n = 10.8625moles

PV = nRT

V    = nRT/P

       = 10.8625*0.0821*296/0.95

       = 277.87L

mass of gases mixture = 100g

The density of gases mixture = mass/volume in L

                                                 = 100/277.87   = 0.36g/L >>>>answer